# Capillarity in Stationary Random Granular Media: Distribution-Aware Screening and Quantitative Supercell Sizing

### v2

@@ II.6 Periodic cell problem for capillarity (screened/modified Helmholtz)
and the linearized equilibrium equation takes the modified (screened) Helmholtz form $\nabla^{2}h-\lambda^{-2}h=0$ in homogeneous media [Kralchevsky2000, Vella2005]. Here, $\rho_{\mathrm{nw}}$ and $\rho_{\mathrm{w}}$ denote the nonwetting and wetting fluid densities, respectively, and $\gamma_{\mathrm{nw}}$ is the interfacial tension of that fluid pair. Thus $\lambda$ is set by the fluid–fluid interface that resides in the pore space. The screening parameters $(\gamma_{\mathrm{nw}},\rho_{\mathrm{w}}-\rho_{\mathrm{nw}})$ pertain to the wetting/nonwetting fluid pair occupying the pore space; the solid skeleton only enters through the geometry (which sets the pore connectivity and $K_{\mathrm{p}}$).
@@
with $h$ is $Y$-periodic, which is a screened Poisson (a.k.a. modified Helmholtz or Yukawa) problem with periodic boundary conditions. The diffusion term spreads $h$, the reaction term relaxes it over the length $\lambda$, and $\rho_{0}$ represents sources/sinks. Because $c(\mathbf{x})\geq 0$ and $K(\mathbf{x})>0$ a.e., the reaction term removes the constant-function kernel that appears in pure diffusion; hence no compatibility condition on $\rho_{0}$ is required. This screened structure and its free-space Green functions are standard [Winkelmann2021].
@@
Periodic finite elements assemble (61) from standard $C^{0}$ bases, while FFT-based solvers leverage periodicity by treating the diffusion part in Fourier space and applying the reaction term pointwise in real space [MoulinecSuquet1998]. A quick units check in the head formulation confirms consistency:
@@
see, e.g., classical results in [BensoussanLionsPapanicolaou2011, JikovKozlovOleinik1994]. For an isotropic scalar summary one may define an apparent capillary length
@@
This formulation is useful for two reasons. First, it provides a numerically robust periodic cell problem that directly returns $K_{\rm app}$ and $\beta_{\rm app}$ from the microstructure, giving a compact macro model (63). Second, the effective decay length $\lambda_{\rm app}$ supplies a physics-based length scale for sizing the periodic supercell in capillarity-dominated problems: the shortest cell edge should be several $\lambda_{\rm app}$ to resolve the long-wavelength (low-$k$) regime of the screened response. Finally, there is a natural connection to pore-morphology approaches: recent work by Alonso–Marroquín and Andersson relates capillary pressure–saturation curves to morphology with minimal physics inputs (contact angle and interfacial tension) [AlonsoAndersson2025PRE]. While their focus is quasistatic drainage/imbibition, our $\lambda_{\rm app}$ plays a complementary role as the dynamic screening length of capillary disturbances in continuum models; it can be compared with pore-throat statistics that control entry pressures in morphology-based analyses.
@@ II.7 Choosing the supercell size for capillarity
Writing the integral range as $A=c_{A}\,\xi^{3}$ with a dimensionless shape factor $c_{A}$ determined by the normalized covariance (e.g., $c_{A}=\pi^{3/2}$ for Gaussian and $c_{A}=8\pi$ for exponential decay), averaging $n$ independent realizations and imposing a coefficient-of-variation target $\mathrm{CV}\leq\varepsilon$ gives the explicit volume rule
@@ III.1 Capillarity redefined through a distribution-aware weighting
The decomposition $C\approx\int C^{(a)}\,da$ neglects cross-size covariances; include $\int\!\!\int W_{\lambda}\,\widehat{C}^{(a,a^{\prime})}\,da\,da^{\prime}$ if strong size–size correlations are present.
@@
To retain the same units as the integral range $A$ (volume), we weight $C$ with the dimensionless kernel $S_{\lambda}(\mathbf{r})=e^{-|\mathbf{r}|/\lambda}$ (whose integral is $8\pi\lambda^{3}$), rather than the normalized Yukawa Green’s function. Define the screened integral range
@@
Thus $A_{\lambda}\to A$ as $\lambda\to\infty$ and $A_{\lambda}=O(\lambda^{3})$ for short-range $C$, preserving the variance law’s units. With $(\phi_{\lambda},A_{\lambda})$ the distribution-aware variance model for the surrogate becomes
@@ III.2 Impact on choosing the supercell size for capillarity
Changing the grain-size distribution thus changes $\xi$ and $a_{\max}$ via two-point statistics and size tails, modifies $K_{\rm app}$ and hence mathbfLmathbf{L}, and updates the precision term through $(\phi_{\lambda},A_{\lambda})$. In practice, estimate $(\xi,A)$ and $(\xi_{\lambda},A_{\lambda})$ from data, extract $a_{\max}$ from a high quantile of sizes, compute $K_{\rm app}$ and $\beta_{\rm app}$ (or $\beta_{\rm app}^{(\lambda)}$) on small cells to obtain mathbfLmathbf{L}, enforce (108), then apply (110) (or (111)–(112)) for the desired $(n,\varepsilon)$, and verify a posteriori that $k_{\min}\ll k_{c}$ and $k_{\min}\ll 1/\Lambda$ as in §II.3.
@@ V Conclusions
Our analysis assumes small-slope capillarity, stationarity with integrable covariance, and first-order periodic homogenization. When the covariance is non-integrable (algebraic long-range correlations), the integral range $A$ diverges and only spectral criteria based on $k_{\min}$ and empirical variance estimates remain predictive. In hyperuniform media with $\widehat{C}(\mathbf{0})=0$ one has $A=0$ and $\mathrm{Var}[P(Y)]$ decays faster than $1/V$; our formulas then overestimate fluctuations and provide conservative lower bounds on the box size. For large interfacial slopes or strongly heterogeneous $\lambda(\mathbf{x})$ (contact-angle variations, pore-scale contrasts), the screened-Helmholtz surrogate is only a coarse proxy, and the sizing rules should be viewed as guidance for choosing starting boxes rather than sharp criteria.

### v3

@@ II.6 Periodic cell problem for capillarity (screened/modified Helmholtz)
and the linearized equilibrium equation takes the modified (screened) Helmholtz form $\nabla^{2}h-\lambda^{-2}h=0$ in homogeneous media [15, 24]. Here, $\rho_{\mathrm{nw}}$ and $\rho_{\mathrm{w}}$ denote the nonwetting and wetting fluid densities, respectively, and $\gamma_{\mathrm{nw}}$ is the interfacial tension of that fluid pair. Thus $\lambda$ is set by the fluid–fluid interface that resides in the pore space. The screening parameters $(\gamma_{\mathrm{nw}},\rho_{\mathrm{w}}-\rho_{\mathrm{nw}})$ pertain to the wetting/nonwetting fluid pair occupying the pore space; the solid skeleton only enters through the geometry (which sets the pore connectivity and $K_{\mathrm{p}}$).
@@
with $h$ is $Y$-periodic, which is a screened Poisson (a.k.a. modified Helmholtz or Yukawa) problem with periodic boundary conditions. The diffusion term spreads $h$, the reaction term relaxes it over the length $\lambda$, and $\rho_{0}$ represents sources/sinks. Because $c(\mathbf{x})\geq 0$ and $K(\mathbf{x})>0$ a.e., the reaction term removes the constant-function kernel that appears in pure diffusion; hence no compatibility condition on $\rho_{0}$ is required. This screened structure and its free-space Green functions are standard [27].
@@
Periodic finite elements assemble (61) from standard $C^{0}$ bases, while FFT-based solvers leverage periodicity by treating the diffusion part in Fourier space and applying the reaction term pointwise in real space [19]. A quick units check in the head formulation confirms consistency:
@@
see, e.g., classical results in [3, 12]. For an isotropic scalar summary one may define an apparent capillary length
@@
This formulation is useful for two reasons. First, it provides a numerically robust periodic cell problem that directly returns $K_{\rm app}$ and $\beta_{\rm app}$ from the microstructure, giving a compact macro model (63). Second, the effective decay length $\lambda_{\rm app}$ supplies a physics-based length scale for sizing the periodic supercell in capillarity-dominated problems: the shortest cell edge should be several $\lambda_{\rm app}$ to resolve the long-wavelength (low-$k$) regime of the screened response. Finally, there is a natural connection to pore-morphology approaches: recent work by Alonso–Marroquín and Andersson relates capillary pressure–saturation curves to morphology with minimal physics inputs (contact angle and interfacial tension) [2]. While their focus is quasistatic drainage/imbibition, our $\lambda_{\rm app}$ plays a complementary role as the dynamic screening length of capillary disturbances in continuum models; it can be compared with pore-throat statistics that control entry pressures in morphology-based analyses.
@@ II.7 Choosing the supercell size for capillarity
Equation (80) also makes clear that the statistical prefactor grows quadratically with the phase contrast $|K_{g}-K_{m}|$. Hence extremely high-contrast pore/solid systems require either larger cells or more realizations to achieve the same coefficient-of-variation target. Numerically, such limits can also degrade the conditioning of FE/FFT solvers if one phase is treated as strictly nonconducting. In practice it is preferable to regularize the vanishing phase by a small floor value $K_{\min}>0$, solve the periodic problem with preserved ellipticity, and verify convergence of observables as $K_{\min}\to 0$. The resulting sizing rule is unchanged in form but becomes increasingly conservative as the contrast grows. Writing the integral range as $A=c_{A}\,\xi^{3}$ with a dimensionless shape factor $c_{A}$ determined by the normalized covariance (e.g., $c_{A}=\pi^{3/2}$ for Gaussian and $c_{A}=8\pi$ for exponential decay), averaging $n$ independent realizations and imposing a coefficient-of-variation target $\mathrm{CV}\leq\varepsilon$ gives the explicit volume rule
@@ III.1 Capillarity redefined through a distribution-aware weighting
The decomposition $C\approx\int C^{(a)}\,da$ is a size-diagonal approximation that neglects cross-size covariances. This is reasonable for weakly correlated polydispersity, but in dense, jammed, or otherwise strongly structured packings one should retain the full cross-covariance contribution
$\iint\widehat{C}^{(a,a^{\prime})}(\mathbf{k})\,W_{\lambda}(\mathbf{k})\,da\,da^{\prime}$
when constructing the distribution-aware weights. Thus the weighting framework does not rely on independent marks; rather, the diagonal form is a simplifying closure.
@@
To retain the same units as the integral range $A$ (volume), we introduce a screened proxy kernel
$S_{\lambda}(\mathbf{r})=e^{-|\mathbf{r}|/\lambda},$
which is dimensionless and whose integral is $8\pi\lambda^{3}$. This kernel is not the exact finite-window variance kernel of the screened problem; rather, it is a low-pass surrogate that captures the dominant attenuation of modes with $k\gtrsim\lambda^{-1}$ while preserving the dimensional structure of the classical integral range. Define the screened integral range
@@
Accordingly, $A_{\lambda}$ should be interpreted as a screened analogue of $A$ that is asymptotically faithful at the level of scaling and ordering of microstructures, not as an exact replacement for the full windowed variance operator. The approximation is most reliable when $C(\mathbf{r})$ is short-ranged and smooth on the scale $\lambda$, so that the same low-$k$ sector dominates both the exact variance and the screened proxy. For sharply oscillatory, strongly anisotropic, or strongly cross-correlated media, the exact spectral-window expression is preferable. Thus $A_{\lambda}\to A$ as $\lambda\to\infty$ and $A_{\lambda}=O(\lambda^{3})$ for short-range $C$, preserving the variance law’s units. With $(\phi_{\lambda},A_{\lambda})$ the distribution-aware variance model for the surrogate becomes
@@ III.2 Impact on choosing the supercell size for capillarity
Changing the grain-size distribution thus changes $\xi$ and $a_{\max}$ via two-point statistics and size tails, modifies $K_{\rm app}$ and hence $\lambda_{\rm app}$, and updates the precision term through $(\phi_{\lambda},A_{\lambda})$. In practice, estimate $(\xi,A)$ and $(\xi_{\lambda},A_{\lambda})$ from data, extract $a_{\max}$ from a high quantile of sizes, compute $K_{\rm app}$ and $\beta_{\rm app}$ (or $\beta_{\rm app}^{(\lambda)}$) on small cells to obtain $L$, enforce (108), then apply (110) (or (111)–(112)) for the desired $(n,\varepsilon)$, and verify a posteriori that $k_{\min}\ll k_{c}$ and $k_{\min}\ll 1/\Lambda$ as in §II.3.
@@ V Conclusions
Our analysis assumes small-slope capillarity, stationarity with integrable covariance, and first-order periodic homogenization. When the covariance is non-integrable (algebraic long-range correlations), the integral range $A$ diverges and only spectral criteria based on $k_{\min}$ and empirical variance estimates remain predictive. In hyperuniform media with $\widehat{C}(\mathbf{0})=0$ one has $A=0$ and $\mathrm{Var}[P(Y)]$ decays faster than $1/V$; our formulas then overestimate fluctuations and provide conservative lower bounds on the box size. For large interfacial slopes, pore-scale contact-line hysteresis, or strongly heterogeneous $\lambda(\mathbf{x})$ (contact-angle variations, local curvature singularities, pore-throat entry events), the screened-Helmholtz surrogate is only a coarse proxy. Accordingly, the present RVE analysis applies to the macroscopic screened capillary field and its finite-size fluctuations, not to strongly nonlinear localized instabilities such as Haines jumps, burst events, or snap-off. In those regimes, the box sizes proposed here should be interpreted as starting scales for coarse-grained representativity, to be complemented by dedicated pore-scale or transient analysis if local instability statistics are the primary observable.